Can this be simplified?
9/4 (x+ 6/5)-11

Answer by stanbon(48540)

9/4 (x+ 6/5)-11
=(9/4)x+(54/20)-220/20
=(9/4)x-166/20
=(9/4)x-(83)/10
There are other "so called" simplified forms
of this problem. You have to decide when
to stop.
Cheers,
Stan H.